A total of 150 participants are selected for a study of risk factors for cardiovascular disease. At baseline (study start), 24 are classified as hypertensive. At 1 year, an additional 12 have developed hypertension, and at 2 years another 8 have developed hypertension. What is the prevalence of hypertension at 2 years in the study? What is the 2 year cumulative incidence of hypertensions?
Answer #1

Period prevalence refers to prevalence measured over an interval of time. It is the proportion of persons with a particular disease or attribute at any time during the interval.

Prevalence of hypertension at 2 years in study = (24 + 12 + 8) / 150 = 44/150 = 0.2933 i.e. 29.33%

2 year cumulative incidence of hypertensions = 24 + 12 + 8 = 44